OpenStudy (anonymous):

A marching band has 14 marchers in the first row, 16 in the second row, 18 in the third row, and so on, for 15 rows. How many marchers are in the last row?

OpenStudy (anonymous):

find the 15th term

OpenStudy (anonymous):

even numbers.. keep adding 2, 15 times

OpenStudy (anonymous):

row 1, N=14 row 2, N=14+2 = 14 + 2*1 row 3, N=14+2+2 = 14+2*2 row 4, N=14+2+2+2 = 14+2*3 and so on for 15th row..?

OpenStudy (anonymous):

Want the formula?

OpenStudy (anonymous):

Stay in school kid , this is a easy and fun problem , i loved these when i had em .

OpenStudy (anonymous):

These are so complicating man I need help

OpenStudy (anonymous):

do you see a trend the numbers I showed above?

OpenStudy (anonymous):

Yea but theres a faster way to do it though

OpenStudy (anonymous):

just keep adding 2 to 14 , 15 times , not a lot of work .

OpenStudy (anonymous):

14 16 18 20 22 24 26 28 30 32 34 36 38 40 42 44 46 48 50

OpenStudy (anonymous):

42?

OpenStudy (anonymous):

Well not 15 times , 14 times . But the answers 42 .

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@cashybabiie

OpenStudy (anonymous):

yes. thats what I meant, \[a_n=a_0+(n-1)d\\ a_0=14\\ n=15\\ d=2\\ \text{find }a_n \]